I NEED HELP IMMEDIATELY PLEASE!!! I AM STUCK!
For the equation in question 1, what value of x would make the equation undefined?
\c=6350/425-3x

Found 2 solutions by Theo, tinbar:
Answer by Theo(13342)

if your equation is:
y = 6350/(425-3x), then the equation would be undefined when (425-3x) = 0
set 425-3x = 0 and solve for x.
add 3x to both sides of that equation to get:
425 = 3x
divide both sides of that equation by 3 to get:
x = 425/3 = 141 and 2/3.
when x = 141 and 2/3, the equation of y = 6350/(425-3x) becomes:
y = 6350 / 0 which is undefined.

Answer by tinbar(133)

When you have a rational function, it becomes undefined usually when there's a divide by 0 error. You cannot divide a number by 0, it doesn't make sense, it is undefined. For instance, we know 6/3 =2, and this is true if and only if 3*2 = 6
So let's try to divide 1/0 and say the result is x. According to the analogy of 6/3=2, we have 1/0=x, and since 3*2=6, we must have x*0=1, but that makes no sense! Anything times 0 is 0, so this value x cannot exist, it is not defined.
So in your question, what value of x, makes the bottom 0? Whatever it is, we need to reject it since that number asks the function to divide by 0, which cannot be done. So you want to solve 425-3x = 0 for x and that'll be the number you want to leave out, or in other words that's the value of x that makes the function undefined